# Andreev bound states in Josephson junctions of semi-Dirac semimetals

### 31 Dec 2023 v1

@@ III Josephson junction
From the four components of the BdG wavefunction, we get $2\times 4=8$ linear homogeneous equations in the $8$ unknown variables $\left(a_{l},\,b_{l},\,c_{l},\,d_{l},\,a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$, which constitute the coefficients of the piecewise-defined wavefunction. In the resulting equations, the overall $x$-independent factors of $e^{i\,K_{z}\,z}$ cancel out. Let $\mathcal{M}$ denote the $8\times 8$ matrix constructed out of the coefficients of the 8 variables. For the equations to be consistent, we need to impose the condition $\text{det}\,\mathcal{M}=0$. This helps us determine the energy eigenvalues of the subgap ABSs, which are localized near the junction since they decay exponentially with the distance from the weak link into the superconducting region on either side.
IV Results
In order to deal with the determinant of a lower dimensional matrix, we first eliminate four of the eight unknown variables by using four of the linear homogeouns equations. Specifically, in our calculations, we first solve for $\left(a_{l},\,b_{l},\,c_{l},\,d_{l}\right)$ in terms of $\left(a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$, using $\psi_{L}(0,z,K_{z})=\psi_{R}(0,z,K_{z})$. These solutions are plugged in the constraint $\partial_{x}\psi_{L}(x,z,K_{z})\big{|}_{x=0}-\partial_{x}\psi_{R}(x,z,K_{z})\big{|}_{x=0}=V_{0}\,\psi_{L}(0,z,K_{z})$, such that $\left(a_{l},\,b_{l},\,c_{l},\,d_{l}\right)$ no longer appear in the equations resulting from it. Now we construct the $4\times 4$ matrix $\tilde{\mathcal{M}}$ out of the coefficients of the four variables, viz., $\left(a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$. The consistency condition now reduces to the vanishing of the matrix $\text{det}\,\tilde{\mathcal{M}}=0$. The equation resulting from this determinant gives a lower order polynomial equation in $e^{i\,\beta}$ (compared to the one obtained from $\text{det}\,\mathcal{M}=0$) and, hence, is easier to solve. In fact, we obtain a quartic equation in the variable $e^{2\,i\,\beta}$.
Employing the steps outlined above, the equation we need to solve reduces to
$z^{4}\left(K_{z}^{2}-25\right)^{2}\left(V_{0}-2\,{\mathcal{K}}_{x}\right)^{2}\left(V_{0}^{2}+4\,{\mathcal{K}}_{x}^{2}\right)-8\,z^{3}\left(K_{z}-5\right)\left(K_{z}+5\right){\mathcal{K}}_{x}^{2}\left(\cos\phi-1\right)\left(2{\mathcal{K}}_{x}-V_{0}\right)\left[2\left(K_{z}^{2}-25\right){\mathcal{K}}_{x}-V_{0}\,K_{z}^{2}\right]$
$+16\,z^{2}\,{\mathcal{K}}_{x}^{2}\sin^{2}\Big{(}\frac{\phi}{2}\Big{)}\left[V_{0}^{2}\,K_{z}^{2}\left(K_{z}^{2}-25\right)-2\,V_{0}\left(2\,K_{z}^{4}-75\,K_{z}^{2}+625\right){\mathcal{K}}_{x}+\left\{4\left(K_{z}^{2}-50\right)K_{z}^{2}+3750\right\}{\mathcal{K}}_{x}^{2}-1250\,{\mathcal{K}}_{x}^{2}\cos\phi\right]$
$+40000\,{\mathcal{K}}_{x}^{4}\sin^{4}\Big{(}\frac{\phi}{2}\Big{)}+80000\,z\,{\mathcal{K}}_{x}^{4}\sin^{4}\left(\frac{\phi}{2}\right)=0\,,$ (24)
where $z=1+e^{2\,i\,\beta}$. The order of the polynomial equation (whose roots we need to determine) is the same as what is found for the case of the linearly dispersing Rarita-Schwinger-Weyl (RSW) semimetal, featuring four bands (rather than two), which was analysed in Ref. [13].
The coefficients of various powers of $z$ in Eq. (IV) are all real and hence we can analyze the nature of the roots by applying the general criteria applicable for a real quartic polynomial equation. First we obtain the associated depressed quartic, which takes the form $z^{4}+q\,z^{2}+\rho\,z+\gamma=0$. Let $d$ be the discriminant. For a given set of values for $E_{F}$, $V_{0}$, and $K_{z}$, we compute numerically the values of $d$, $q$, and $\gamma$. For each case, we find that $d>0$ for admissible values of $E_{F}$, $V_{0}$, and $K_{z}$. This means that if $q>0$ or $04\,\gamma-q^{2}>$ 0, we get a pair of complex conjugate roots. This is what we find from our numerical simulations. However, since $z=1+e^{2\,i\,\beta}$, as an admissible solution must satisfy the condition $|z-1|=1$. From our numerical data, we find that only one of the pair of complex conjugate roots fall in the allowed category. This is in contrast with the RSW case [13], where we get, in general, four distinct solutions for $|\varepsilon|$.
From the above analysis, we arrive at the conclusion that the energies of the subgap states appear as the pairs ±$\pm|\varepsilon|$ for a doubly degenerate value of $|\varepsilon|$. Therefore, for each value of $|\varepsilon|$, we get a total of four ABSs — two with the value $|\varepsilon|$ and two with the value $-|\varepsilon|$. To demonstrate the results, we include some representative plots, and all of these are obtained by setting $E_{F}=1$. In Fig. 2, we show the behaviour of $|\varepsilon|$ as a function of (a) $K_{z}$ (with a fixed value of $\phi$), and (b) $\phi$ (with a fixed value of $K_{z}$), for some representative values of $V_{0}$. The bound state energies are periodic in $\phi$ with period $2\pi$. They are, in fact, functions of $\cos\phi$, as is evident from Eq. (IV). This dependence is reflected in the nature of the curves in Fig. 2(b). Fig. 3(a) illustrates the variation of $|\varepsilon|$-values against the $\phi\,$-$K_{z}$ plane and, hence, shows the dependence of the bound state energies on both these variables in a combined way.
@@ IV Results
Figure 3: (a) Magnitude of the energy $\varepsilon$ of the Andreev bound states against the $\phi\,$-$K_{z}$ plane for $E_{F}=1$ and $V_{0}=100$. (b) The behaviour of the total Josephson current ($\propto\tilde{I}$), in arbitrary units, as a function of $\phi$, obtained at $E_{F}=1$ and $k_{B}\,T=0.005\,\Delta_{0}$. We have used four values of $V_{0}$ as shown in the plotlegends.
@@
where $\varepsilon_{n}$ labels the energy values of the four Andreev bound states and $f(\zeta)=1/\left(1+e^{\frac{\zeta}{k_{B}\,T}}\right)$ is the Fermi-Dirac distribution. Fig. 3(b) shows the behaviour of $I_{J}$ as a function of $\phi$, scaled by appropriate numbers/variables (this scaled quantity being denoted as $\tilde{I}$), for four values of $V_{0}$.
V Summary and outlook
@@
In this paper, we have have computed the characteristics of the emergent ABSs, and the resulting Josephson current, in a Josephson junction built with a two-dimensional semi-Dirac semimetal. The junction is taken to be perpendicular to the quadratically dispersing direction in this anisotropic material, featuring hybrid dispersion. The weak link between the two superconducting phases is modelled by a Dirac delta function potential. Using the BdG formalism, we have determined the wavefunction localizing at the junction for $|\varepsilon|\ll\Delta_{0}$. This requires a piecewise continuous definition — the usual procedure for solving reflection and transmission problems in quantum mechanics involving junction configurations. Demanding consistency of the equations obtained from matching the boundary conditions at the junction, we arrive at a quartic polynomial equation in an appropriate variable variable (related to the modulus of the ABS energy $\varepsilon$), obtained from the vanishing of the relevant determinant. The physically admissible roots of this quartic give the discrete energy spectrum $\varepsilon$ of the ABSs. Although a closed form solution cannot be found due to the fourth order polynomial involved, we have deduced the nature of the roots from our numerical data. We have found that the values of $|\varepsilon|$ are doubly degenerate, leading to four ABSs — two with energy $|\varepsilon|$ and two with energy $-|\varepsilon|$. We have also shown that the solutions depend on the phase difference ($\phi$) between the two superconducting regions via functions of $\cos\phi$.
@@
Our main finding is that, because of the quadratic dispersion, we need to include the “evanescent” wave solutions while defining the wavefunctions in each region. This is stark with the cases where the propagation axis is along a direction of linear-in-momentum dispersion [5, 8, 10, 11, 12, 13]. The existence of the extra solutions results in a higher order polynomial equation to be solved, compared to the analogous linearly dispersing cases [5, 8, 10, 11, 12] with the same number of bands involved. In fact, comparing our results with the isotropic four-band RSW system studied earlier [13], the order of the polynomial for propagation along the quadratic-in-momentum dispersion in a two-band system is the same as that in a linearly dispersing four-band system.
@@
In future, it will be worthwhile to study the Josephson effect in 3d multi-Weyl semimetals [11, 12, 39, 40], by considering the Josephson junction aligned perpendicular to one of the nonlinearly dispersing directions. Another interesting avenue is to introduce a tilt in the Hamiltonian [12] and investigate the resulting ABSs. Last but not the least, analysis of Josephson junctions built with isotropic Luttinger semimetals harbouring quadratic band crossing points [42, 43] is left for future work.

### 08 Apr 2024 v2

@@ 3 Josephson junction
From the four components of the BdG wavefunction, we get $2\times 4=8$ linear homogeneous equations in the $8$ unknown variables $\left(a_{l},\,b_{l},\,c_{l},\,d_{l},\,a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$, which constitute the coefficients of the piecewise-defined wavefunction. In the resulting equations, the overall $z$-independent factors of $e^{i\,K_{z}\,z}$ cancel out. Let $\mathcal{M}$ denote the $8\times 8$ matrix constructed out of the coefficients of the 8 variables. For the equations to be consistent, we need to impose the condition $\text{det}\,\mathcal{M}=0$. This helps us determine the energy eigenvalues of the subgap ABSs, which are localized near the junction, since they decay exponentially with the distance from the weak link into the superconducting region on either side.
4 Results
@@
Figure 3: (a) Magnitude of the energy $\varepsilon$ of the Andreev bound states against the $\phi\,$-$K_{z}$ plane, for $E_{F}=1$ and $V_{0}=100$. (b) The behaviour of the total Josephson current ($\propto\tilde{I}$), in arbitrary units, as a function of $\phi$, obtained at $E_{F}=1$ and $k_{B}\,T=0.005\,\Delta_{0}$. We have used four values of $V_{0}$ as shown in the plotlegends.
In order to reduce the complexity of the computations to the determinant of a lower dimensional matrix, we first eliminate four of the eight unknown variables by using four of the eight linear homogeneous equations. Specifically, in our calculations, we first solve for $\left(a_{l},\,b_{l},\,c_{l},\,d_{l}\right)$ in terms of $\left(a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$, using $\psi_{L}(0,z,K_{z})=\psi_{R}(0,z,K_{z})$. These solutions are plugged into the constraint
$\partial_{x}\psi_{R}(x,z,K_{z})\big{|}_{x=0}-\partial_{x}\psi_{L}(x,z,K_{z})\big{|}_{x=0}=V_{0}\,\psi_{L}(0,z,K_{z})\,,$
such that $\left(a_{l},\,b_{l},\,c_{l},\,d_{l}\right)$ no longer appear in the equations resulting from the operation. Now we construct the $4\times 4$ matrix $\tilde{\mathcal{M}}$ out of the coefficients of the four variables, viz., $\left(a_{r},\,b_{r},\,c_{r},\,d_{r}\right)$. The explicit form of this matrix is very long and, therefore, we write the 4 columns one by one:
$c_{1}=\begin{bmatrix}e^{i\,\beta}\left(E_{F}^{2}-K_{z}^{2}\right)\left[e^{i\,\phi}\left\{E_{F}\left(-z\,V_{0}+(1+i)(1+\left(1+i\right)z)\kappa_{x}\right)+K_{z}\left(-z\,V_{0}+2\,i\left(z+1\right)\kappa_{x}\right)\right\}-\left(1+i\right)\kappa_{x}\left\{E_{F}+\left(1+i\right)K_{z}\right\}\right]\\ \\ e^{i\,\beta}\kappa_{x}^{2}\left(E_{F}+K_{z}\right)\left[e^{i\,\phi}\left\{E_{F}\left(-z\,V_{0}+\left(1+i\right)(1+\left(1+i\right)z)\kappa_{x}\right)+K_{z}\left(z\,V_{0}-2\,i\left(z+1\right)\kappa_{x}\right)\right\}-\left(1+i\right)E_{F}\,\kappa_{x}+2\,i\,K_{z}\,\kappa_{x}\right]\\ \\ \left(E_{F}^{2}-K_{z}^{2}\right)\left[E_{F}\left\{-z\,V_{0}+\left(1+i\right)\kappa_{x}\left(\left(z+1\right)e^{i\,\phi}+i\,z-1\right)\right\}+K_{z}\left\{-z\,V_{0}+2\,i\left(-1+e^{i\,(2\,\beta+\phi)}\right)\kappa_{x}\right\}\right]\\ \\ \kappa_{x}^{2}\left(E_{F}+K_{z}\right)\left[E_{F}\left\{-z\,V_{0}+\left(1+i\right)\kappa_{x}\left(\left(z+1\right)e^{i\,\phi}+i\,z-1\right)\right\}+K_{z}\left\{z\,V_{0}-2\,i\left(-1+e^{i\,(2\,\beta+\phi)}\right)\kappa_{x}\right\}\right]\end{bmatrix},$ (33)
$c_{2}=\begin{bmatrix}\left(E_{F}^{2}-K_{z}^{2}\right)\left[e^{i\,\phi}\left\{V_{0}\,z\left(E_{F}-K_{z}\right)+\left(1+i\right)E_{F}\left(z-i\right)\kappa_{x}+2\,i\,K_{z}\,\kappa_{x}\right\}-\left(1-i\right)\left(z+1\right)\kappa_{x}\left\{E_{F}-\left(1-i\right)K_{z}\right\}\right]\\ \\ \kappa_{x}^{2}\left(E_{F}-K_{z}\right)\left[e^{i\,\phi}\left\{-V_{0}\,z\left(E_{F}+K_{z}\right)+\left(1+i\right)\kappa_{x}\left(\left(1+i\right)K_{z}-E_{F}\left(z-i\right)\right)\right\}+\left(1-i\right)\left(z+1\right)\kappa_{x}\left\{E_{F}+\left(1-i\right)K_{z}\right\}\right]\\ \\ e^{i\,\beta}\left(E_{F}^{2}-K_{z}^{2}\right)\left[E_{F}\left\{V_{0}\,z+\left(1+i\right)\kappa_{x}\left(\left(1+i\right)z-i\,e^{i\,\phi}+i\right)\right\}+K_{z}\left\{-V_{0}\,z+2\,i\,\kappa_{x}\left(e^{i\,\phi}-z-1\right)\right\}\right]\\ \\ e^{i\,\beta}\kappa_{x}^{2}\left(E_{F}-K_{z}\right)\left[K_{z}\left\{-V_{0}\,z+2\,i\,\kappa_{x}\left(e^{i\,\phi}-z-1\right)\right\}-E_{F}\left\{V_{0}\,z+\left(1+i\right)\kappa_{x}\left(\left(1+i\right)z-i\,e^{i\,\phi}+i\right)\right\}\right]\end{bmatrix},$ (41)
$c_{3}=\begin{bmatrix}e^{i\left(\beta+\phi\right)}\left(E_{F}^{2}-K_{z}^{2}\right)\left[-z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left\{-V_{0}\,z-\left(1-i\right)\kappa_{x}\left(\left(1+i\right)z-e^{-i\,\phi}+1\right)\right\}\right]\\ \\ e^{i\,\beta}\,\kappa_{x}^{2}\left(E_{F}+K_{z}\right)\left[e^{i\,\phi}\left\{-z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left(V_{0}\,z+\left(1-i\right)\kappa_{x}+2\,z\,\kappa_{x}\right)\right\}-\left(1-i\right)E_{F}\,\kappa_{x}\right]\\ \\ \left(E_{F}^{2}-K_{z}^{2}\right)\left[-z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)-E_{F}\left\{V_{0}\,z+\left(1+i\right)\kappa_{x}\left(-i\,e^{i\,\phi}\left(z+1\right)+z+i\right)\right\}\right]\\ \\ \kappa_{x}^{2}\left(E_{F}+K_{z}\right)\left[-z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left\{V_{0}\,z+\left(1+i\right)\kappa_{x}\left(-i\,e^{i\,\phi}\left(z+1\right)+z+i\right)\right\}\right]\end{bmatrix},$ (49)
$c_{4}=\begin{bmatrix}\left(E_{F}^{2}-K_{z}^{2}\right)\left[e^{i\,\phi}\left\{-z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left(V_{0}\,z+\left(1-i\right)\left(z-i\right)\kappa_{x}\right)\right\}+\left(1+i\right)E_{F}\left(z+1\right)\kappa_{x}\right]\\ \\ \kappa_{x}^{2}\left(E_{F}-K_{z}\right)\left[e^{i\,\phi}\left\{z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left(V_{0}\,z+\left(1-i\right)\left(z-i\right)\kappa_{x}\right)\right\}+\left(1+i\right)\left(z+1\right)E_{F}\,\kappa_{x}\right]\\ \\ -e^{i\,\beta}\left(E_{F}^{2}-K_{z}^{2}\right)\left[-z\,E_{F}\,V_{0}+\left(1+i\right)E_{F}\,\kappa_{x}\left\{\left(i-1\right)z+e^{i\,\phi}-1\right\}+z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)\right]\\ \\ e^{i\,\beta}\kappa_{x}^{2}\left(E_{F}-K_{z}\right)\left[z\,K_{z}\left(V_{0}+2\,\kappa_{x}\right)+E_{F}\left\{V_{0}\,z-\left(1+i\right)\kappa_{x}\left(\left(i-1\right)z+e^{i\,\phi}-1\right)\right\}\right]\end{bmatrix},$ (57)
such that
~$\tilde{\mathcal{M}}=e^{-i\,\phi}\,\begin{bmatrix}c_{1}\quad c_{2}\quad c_{3}\quad c_{4}\end{bmatrix}\,.$
Here, $z=e^{2\,i\,\beta}-1$. The consistency condition reduces to $\text{det}\,\tilde{\mathcal{M}}=0$. The equation resulting from this vanishing determinant gives a lower order polynomial equation in $e^{i\,\beta}$ (compared to the one obtained from $\text{det}\,\mathcal{M}=0$) and, hence, is easier to solve. In fact, we obtain a quartic equation in the variable $z$, as shown below:
$z^{4}\left(K_{z}^{2}-E_{F}^{2}\right)^{2}\left(V_{0}+2\,{\mathcal{K}}_{x}\right)^{2}\left(V_{0}^{2}+4\,{\mathcal{K}}_{x}^{2}\right)-8\,z^{3}\left(K_{z}^{2}-E_{F}^{2}\right){\mathcal{K}}_{x}^{2}\left(\cos\phi-1\right)\left(2\,{\mathcal{K}}_{x}+V_{0}\right)\left[2\left(K_{z}^{2}-E_{F}^{2}\right){\mathcal{K}}_{x}+V_{0}\,K_{z}^{2}\right]$
$+16\,z^{2}\,{\mathcal{K}}_{x}^{2}\,\sin^{2}\Big{(}\frac{\phi}{2}\Big{)}\left[-E_{F}^{2}\,K_{z}^{2}\left(V_{0}+2\,\kappa_{x}\right)\left(V_{0}+4\,\kappa_{x}\right)+2\,E_{F}^{4}\,\kappa_{x}\left(V_{0}+3\,\kappa_{x}\right)+K_{z}^{4}\left(V_{0}+2\,\kappa_{x}\right)^{2}-2\,E_{F}^{4}\,{\mathcal{K}}_{x}^{2}\cos\phi\right]$
$+128\,z\,E_{F}^{2}\,{\mathcal{K}}_{x}^{4}\,\sin^{4}\left(\frac{\phi}{2}\right)+64\,E_{F}^{2}\,{\mathcal{K}}_{x}^{4}\,\sin^{4}\Big{(}\frac{\phi}{2}\Big{)}=0\,.$ (58)
The order of this polynomial equation (whose roots we need to determine) is the same as what we found for the case of the linearly dispersing Rarita-Schwinger-Weyl (RSW) semimetal [14], featuring four bands (rather than two).
The coefficients of various powers of $z$ in Eq. (4) are all real and, hence, we can analyze the nature of the roots by applying the general criteria applicable for a real quartic polynomial equation. First we obtain the associated depressed quartic, which takes the form:
$z^{4}+q\,z^{2}+\rho\,z+\gamma=0\,.$
Let $\mathcal{D}$ be the discriminant of the depressed polynomial. For a given set of values for $E_{F}$, $V_{0}$, and $K_{z}$, we compute numerically the values of $\mathcal{D}$, $q$, and $\gamma$. For each case, we find that $\mathcal{D}>0$. This means that if $q>0$ or $04\,\gamma-q^{2}>$ 0, we get a pair of complex conjugate roots (i.e., we get four complex roots which are of the form $z_{1}$, $z_{1}^{*}$, $z_{2}$, and $z_{2}^{*}$). This is what we find from our numerical simulations. However, since $z=e^{2\,i\,\beta}-1$, an admissible solution must satisfy the condition $|z+1|=1$. From our numerical data, we find that only one set of complex conjugate roots, out of the pair, fall in the allowed category. This is in contrast with the RSW case [14], where we get, in general, four distinct solutions for $|\varepsilon|$.
From the above analysis, we arrive at the conclusion that the energies of the subgap states appear as the pairs ±$\pm|\varepsilon|$ for a doubly degenerate value of $|\varepsilon|$. Therefore, for each value of $|\varepsilon|$, we get a total of four ABSs — two with the value $|\varepsilon|$ and two with the value $-|\varepsilon|$. To demonstrate the results, we include some representative plots, and all of these are obtained by setting $E_{F}=1$. In Fig. 2, we show the behaviour of $|\varepsilon|$ as a function of (a) $K_{z}$ (with a fixed value of $\phi$), and (b) $\phi$ (with a fixed value of $K_{z}$), for some representative values of $V_{0}$. The bound state energies are periodic in $\phi$ with period $2\pi$. They are, in fact, functions of $\cos\phi$, as is evident from Eq. (4). This dependence is reflected in the nature of the curves in Fig. 2(b). Fig. 3(a) illustrates the variation of $|\varepsilon|$-values against the $\phi\,$-$K_{z}$ plane and, hence, shows the dependence of the bound state energies on both these variables in a combined way.
@@
where $\varepsilon_{n}$ labels the energy values of the four Andreev bound states, and $f(\zeta)=1/\left(1+e^{\frac{\zeta}{k_{B}\,T}}\right)$ is the Fermi-Dirac distribution function. Fig. 3(b) shows the behaviour of $I_{J}$ as a function of $\phi$, scaled by appropriate numbers/variables (this scaled quantity being denoted as $\tilde{I}$), for four values of $V_{0}$.
5 Summary and outlook
@@
In this paper, we have have computed the characteristics of the emergent ABSs, and the resulting Josephson current, in a Josephson junction built with a 2d semi-Dirac semimetal. The junction is taken to be perpendicular to the quadratically dispersing direction in this anisotropic material, which features a hybrid dispersion. The weak link between the two superconducting regions is modelled by a Dirac delta function potential. Using the BdG formalism, we have determined the wavefunction localizing at the junction for $|\varepsilon|\ll\Delta_{0}$. This requires a piecewise continuous definition — the usual procedure adopted for solving reflection and transmission problems in quantum mechanics involving junction configurations. Demanding consistency of the equations obtained from matching the boundary conditions at the junction, we arrive at a quartic polynomial equation in an appropriate variable variable (related to the modulus of the ABS energy $\varepsilon$), obtained from the vanishing of the relevant determinant. The physically admissible roots of this quartic give the discrete energy spectrum $\varepsilon$ of the ABSs. Although a closed form solution cannot be found, due to the fourth order polynomial involved, we have deduced the nature of the roots from our numerical data. We have found that the values of $|\varepsilon|$ are doubly degenerate, leading to four ABSs — two with energy $|\varepsilon|$ and two with energy $-|\varepsilon|$. We have also shown that the solutions depend on the phase difference ($\phi$) between the two superconducting regions via functions of $\cos\phi$.
@@
Our main finding is that, because of the quadratic dispersion, we need to include the “evanescent” wave solutions, while defining the wavefunctions in each region. This is in stark contrast with the cases where the propagation axis is along a direction of linear-in-momentum dispersion [5, 9, 11, 12, 13, 14]. The existence of the extra solutions results in a higher order polynomial equation to be solved, compared to the analogous linearly dispersing cases [5, 9, 11, 12, 13] with the same number of bands involved. In fact, comparing our results with the isotropic four-band RSW system studied earlier [14], we find that the order of the polynomial, for propagation along the quadratic-in-momentum dispersion in a two-band system, is the same as that in a linearly dispersing four-band system.
In this paper, we have considered the Josephson current in the thin-barrier-limit, which allows one to model the barrier as a Dirac delta function. One can consider a finite barrier instead, but the computations will be significantly challenging. In such cases, we have a finite normal state region sandwiched between two superconducting regions, and we need to determine the piecewise continuous wavefunctions from the boundary-condition-matching at the S-B and B-S junctions at $x=0$ and $x=\ell$, where $\ell$ is the length of the barrier along the propagation direction. Furthermore, barriers of more generic shapes (rather than the simple rectangular form) can be considered, which might be realized in futuristic experimental set-ups. Yet another interesting barrier-configuration is to consider scenarios where the anisotropic dispersion is rotated about the propagation-axis across the junction(s), as considered in Refs. [44, 45].
@@
In the future, it will be worthwhile to study the Josephson effect in 3d multi-Weyl semimetals [12, 13, 41, 42], by considering the Josephson junction aligned perpendicular to one of the nonlinearly dispersing directions. Another interesting avenue is to introduce a tilt in the Hamiltonians of various nodal-point semimetals [13], and investigate the resulting ABSs. Last but not the least, analysis of Josephson junctions built with isotropic Luttinger semimetals, harbouring quadratic band crossing points [46, 47], is left for future work.
Acknowledgments
We thank Angad Singh for a careful reading of the manuscript.